OpenStudy (anonymous):

Express in terms of i. 4sqrt-225

OpenStudy (lgbasallote):

\[4\sqrt{-225}\] ???

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

here's a hint: \[\huge 15^2 = 225\] does that help?

OpenStudy (anonymous):

not really

OpenStudy (unklerhaukus):

\[\sqrt {ab}=\sqrt a\sqrt b\]

OpenStudy (lgbasallote):

how about this: \[\huge (15i)^2 \implies (225)(-1) \implies -225\]

OpenStudy (lgbasallote):

do you get what im implying now?

OpenStudy (anonymous):

not really.

OpenStudy (lgbasallote):

.....what it means is that \[\huge \sqrt{-225} = 15i\]

OpenStudy (anonymous):

what about the 4 in front tho

OpenStudy (lgbasallote):

you multiply it to 15i

OpenStudy (lgbasallote):

what's 15*4?

OpenStudy (anonymous):

60

OpenStudy (lgbasallote):

so what do you think will be the final answer?

OpenStudy (anonymous):

\[\sqrt[4]{-255}=15i*4=60\]

OpenStudy (lgbasallote):

where did i go?

OpenStudy (lgbasallote):

wait.... is it \[\huge \sqrt[4]{-255} \; \; \text{or} \; \; 4\sqrt{-225}\] ???

OpenStudy (anonymous):

the second one sorry still learning the things on here

OpenStudy (lgbasallote):

so anyway... \[15i \times 4\] what do you think is the product?

OpenStudy (anonymous):

\[4\sqrt{-255}=15i*4=60i\]

OpenStudy (lgbasallote):

right